OpenStudy (anonymous):

Find the area of a rectangular object which has a length of 5√8 inches and a width of 2√5 inches. Would the answer be 4√10 or 20√10?

jimthompson5910 (jim_thompson5910):

Area = Length * Width A = LW A = 5√8 * 2√5 A = (5*2)*(√8*√5) A = 10√(8*5) A = 10√(4*2*5) A = 10√4*√(2*5) A = 10*2*√(2*5) A = 20√10 So your second choice is correct.